Suppose X1 and X2 are random variables of the discrete type which have the joint pmf p(x1,x2)=(x1+2x2)/18, (x1,x2)=(1,1),(1,2),(2,1),(2,2), zero elsewhere. Determine the conditional mean and variance of X2,given X1=x1, for x1=1 or 2. Also ,compute E(3X1-2X2)


1
Expert's answer
2021-10-17T16:43:32-0400

Let "x_1=1"

Conditional distribution "f(x_2|x_1=1)" is:

"\def\arraystretch{1.5}
 \begin{array}{c:c:c}
 x_2 & 1 & 2 \\ \hline
 f(x_2|x_1=1) & \frac{1/6}{8/18}=\frac{3}{8} &\frac{ 5}{8} \\
 \end{array}"

Then conditional mean

"M(x_2|x_1=1)=1\cdot \frac{3}{8}+2\cdot \frac{5}{8}=\frac{13}{8};"

"M(x_2^2|x_1=1)=1\cdot \frac{3}{8}+4\cdot \frac{5}{8}=\frac{23}{8}"

Conditional variance

"Var(x_2|x_1=1)=M(x_2^2|x_1=1)-\left( M(x_2|x_1=1) \right)^2=\\
\frac{23}{8}-\left(\frac{13}{8}\right)^2=\frac{15}{64}"

Now let be "x_1=2"

In this case Conditional distribution "f(x_2|x_1=2)" is:

"\def\arraystretch{1.5}
 \begin{array}{c:c:c}
 x_2 & 1 & 2 \\ \hline
 f(x_2|x_1=2) & \frac{\frac{4}{18}}{\frac{10}{18}}=\frac{2}{5} & \frac{3}{5} \\
 \end{array}"

And further we have:

"M(x_2|x_1=2)=1\cdot \frac{2}{5}+2\cdot \frac{3}{5}=\frac{8}{5};"

"M(x_2^2|x_1=2)=1\cdot \frac{2}{5}+4\cdot \frac{3}{5}=\frac{14}{5}"

"Var(x_2|x_1=2)=\frac{14}{5}-(\frac{8}{5})^2=\frac{6}{25}"

For computing "E(3X_1-2X_2)" we use the table of "(x_1,x_2)" distribution:

"\def\arraystretch{1.5}
 \begin{array}{c:c:c:c:c}
 (x_1,x_2) & (1,1) & (1,2)&(2,1)&(2,2) \\ \hline
 f(x_1,x_2) & \frac{3}{18} & \frac{5}{18}&\frac{4}{18}&\frac{6}{18}\\
 \hdashline
 3\cdot x_1-2\cdot x_2 & 1 & -1&4&2
\end{array}"

Therefore

"E(3X_1-2X_2)=\frac{3\cdot 1+5\cdot(-1)+4\cdot 4+6\cdot 2}{18}=\frac{26}{18}=\frac{13}{9}"